OpenStudy (anonymous):

the sum of two numbers is 10. if one number is subtracted from the other the result is -4. find two numbers

OpenStudy (anonymous):

let a and b be the numbers then: a+b=10 a-b=-4 solve the equations

OpenStudy (alfie):

x+y = 10 x-y = -4 x=3 y=7